# Learning Domain-Invariant Features for Out-of-Context News Detection

### v1

@@ Abstract.
Multimodal out-of-context news is a common type of misinformation on online media platforms. This involves posting a caption, alongside an invalid out-of-context news image. Reflecting its importance, researchers have developed models to detect such misinformation. However, a common limitation of these models is that they only consider the scenario where pre-labeled data is available for each domain, failing to address the out-of-context news detection on unlabeled domains (e.g., unverified news on new topics or agencies). In this work, we therefore focus on domain adaptive out-of-context news detection. In order to effectively adapt the detection model to unlabeled news topics or agencies, we propose ConDA-TTA (Contrastive Domain Adaptation with Test-Time Adaptation) which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature. In addition, it leverages target domain statistics during test-time to further assist domain adaptation. Experimental results show that our approach outperforms baselines in 5 out of 7 domain adaptation settings on two public datasets, by as much as 2.93% in F1 and 2.08% in accuracy.
contrastive learning, domain adaptation, out-of-context news detection, multimodal learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference XXXX; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/24/10ccs: Computing methodologies Natural language processingccs: Information systems Web applicationsccs: Computing methodologies Computer vision
1. Introduction
@@
Online news platforms suffer from the release and spread of misinformation. Disseminating misinformation is especially concerning in news topics related to key societal issues, such as public health and politics, where it could result in misleading public awareness and arousing public panic. A common (Fazio, 2020) yet subtle type of misinformation is image repurposing, also called out-of-context news. Unlike DeepFake (Dolhansky et al., 2020) models which generate non-existing images, this approach attaches a false claim with a real image outside of its original context (see Figure 1). This lowers the technical threshold (Fazio, 2020) of producing misinformation since anyone can trivially attach a fabricated caption to an image and post it online, thereby enlarging its scope of occurrence. Complicating matters, out-of-context news cannot be detected using a sole modality input (as in isolation, each modality may be valid).
@@
Figure 1. Examples of out-of-context news of three different news topics from the Twitter-COMMs dataset.
@@
Out-of-context news detection has received a growing attention in recent years. Existing works (Biamby et al., 2022; Luo et al., 2021; Abdelnabi et al., 2022; Mu et al., 2023) fine-tune pre-trained vision and text encoders to represent the image and text, and then devise neural networks to measure vision-text consistency. More recent works (Dai et al., 2024; Zhang et al., 2023) query the Multimodal Large Language Model (MLLM) to obtain the prediction with explanations. Nevertheless, their approaches do not address the adaptation to entirely new topics or news agencies. Due to the wide array of news topics shared online, it is infeasible to re-train the detector every time a new topic emerges. Therefore, a detection model that can adapt to new topics will greatly relieve the operational burden. Additionally, considering annotation costs, it is equally critical to build a detection model that is able to adapt to other news agencies with minimal efforts, so that the model developed for one news agency can be easily extended to other popular news agencies, e.g. USA Today. Because of this, we argue that it is important to devise domain adaptive out-of-context news detection model.
@@
Unfortunately, prior works neglect addressing the domain adaption issue on out-of-context news detection. Trained on sufficient labeled data from each domain, existing models tend to capture domain-specific knowledge which is not shared among different domains. Although this knowledge is helpful for improving the model’s performance on annotated domains, it results in inferior performance on unannotated domains (Wang et al., 2018). In light of this, we believe that learning domain-invariant features will help the detection model adapt to unlabeled domains. Though important, enabling domain adaptive capabilities on this task comes with several challenges. First, existing out-of-context news datasets only cover a small number of topics/agencies. This makes it challenging for the detection model to learn domain-invariant features, because the learned features might be biased towards the limited number of source domains. Second, in (social media) news dataset, the data distributions of different domains vary, since news from different domains tend to have different writing styles and different image styles. This makes domain-invariant knowledge learning challenging. It is critical to ensure that transferable domain-invariant knowledge is actually learned. Otherwise, the model may bypass learning transferable domain knowledge and end up learning semantic patterns as a shortcut to fulfill the classification task (Li et al., 2023a; Chi et al., 2023).
@@
To tackle the aforementioned challenges, we propose Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA) for domain adaptive out-of-context news detection. ConDA-TTA first uses an MLLM to directly encode the image and text into a multimodal feature representation. After feature encoding, we adopt the contrastive loss to learn a representation in the projected space where the original news stays further away from the out-of-context news. To overcome the shortage of labeled source domains, we use unlabeled target domain data as well, to learn less biased domain-invariant features. We further apply Maximum Mean Discrepancy (MMD) to measure and reduce the discrepancy of the learned representations between the source and target domain. In this way, MMD ensures that the domain-invariant features are captured from the learned representations. Finally, we incorporate Test-Time Adaptation (TTA) to update the statistic-related model parameters in the evaluation phase so that the model can better adapt to the target domain.
@@
  • To the best of our knowledge, we are the first to investigate domain adaptation in out-of-context news detection. To address the challenges, we propose a novel approach using contrastive domain adaptation with test-time adaptation (ConDA-TTA).
@@
  • We evaluate our approach and demonstrate that it achieves state-of-the-art (SOTA) performance under most domain adaptation settings. Notably, it outperforms the baseline by as much as 2.93% in F1 and 2.08% in accuracy when the domain is defined as news topic. Additionally, it outperforms the baseline by as much as 1.82% in F1 and 1.84% in accuracy when the domain is defined as news agency.
@@
  • We conduct a comprehensive ablation study and show that MMD tends to be the most contributing component to the domain adaptation on the Twitter-COMMs dataset, while TTA tends to be the most contributing component on the NewsCLIPpings dataset.
2. Related Work
2.1. Out-of-Context News Detection
@@
Prior works on out-of-context news detection mainly adopt three approaches: (i) fine-tuning or (ii) prompting large vision and language models, and (iii) leveraging synthetic information. Biamby et al. (2022) fine-tunes CLIP (Radford et al., 2021) and a linear classifier to detect out-of-context news. Similarly, Luo et al. (2021) fine-tunes both CLIP and VisualBert (Li et al., 2019) with a linear classifier. Abdelnabi et al. (2022) concatenates the retrieved textual evidence with visual evidence and then inputs the learned representations into a simple classifier. These works involve fine-tuning a large multimodal pre-trained model (e.g. CLIP), which is very compute-intensive (e.g.  (Biamby et al., 2022) uses 8 days of training on 10 GPUs), and does not enable domain adaptation.
@@
Later work leverages synthetic multimodal information. Shalabi et al. (2023) uses a stable diffusion model (Dhariwal and Nichol, 2021) to generate augmented image from text, and a pre-trained image captioning model to generate augmented text from image. (Yuan et al., 2023) extracts stances from the external multimodal evidences and uses it to enhance evidence-based out-of-context detection. More recent works focus on enhancing the interpretability. (Qi et al., 2024) adopts two-stage instruction tuning on Instruct-BLIP (Dai et al., 2024) to provide accurate and persuasive explanations to the prediction. (Zhang et al., 2023) uses neural symbolic model to enhance model interpretability. (Shalabi et al., 2024) fine-tunes the MiniGPT-4 to detect out-of-context news. Although effective, the above works do not address the domain adaptation problem in out-of-context news detection.
2.2. Domain Adaptive Fake News Detection
@@
To the best of our knowledge, no prior work has explored domain adaptive out-of-context news detection. However, some prior works have investigated techniques to enable domain adaptation in multimodal fake news detection — a task close to out-of-context news detection. Some works (Wang et al., 2018; Zhang et al., 2020, 2021; Yuan et al., 2021) adopt adversarial learning through a domain classifier/discriminator to learn the domain-invariant feature. On text-based domain adaptation in fake news detection, Nan et al. (2022) adopts meta learning and Ding et al. (2022) further combines it with adversarial learning. Furthermore, incorporating user comments and user-news interaction information, Mosallanezhad et al. (2022) adopts reinforcement learning that exploits cross-domain and within-domain knowledge to achieve robustness in the target domain. Yue et al. (2022) proposes a contrastive domain adaptation method in order to reduce the intra-class discrepancy and enlarge the inter-class discrepancy. Lin et al. (2023) employs prompt engineering to learn language-agnostic contextual representations and models the domain-invariant structural features from the propagation threads. In contrast, our work focuses on the domain adaptation in out-of-context news detection and adopts both contrastive learning and test-time adaptation to address the unique challenges (Section  1) of this task.
2.3. Test-Time Adaptation
Test-time adaptation (TTA) is an emerging learning paradigm that increases the trained model’s generalization ability to the test samples under distribution shifts (Liang et al., 2023). We later employ this to improve our performance. Many works focus on adapting batch normalization (BN) layer’s parameters during model testing: Li et al. (2016) updates the BN statistics; Wang et al. (2020) further updates the learnable parameters using information entropy; Zhao et al. (2023) proposes to use exponential moving average (EMA) to update BN parameters and to address the imbalanced classes problem. Wang et al. (2022) proposes continual test-time adaptation for reducing error accumulation and catastrophic forgetting. Recent works propose to conduct sample filtering during TTA: Niu et al. (2022) filters out less reliable samples and maintains sample diversity at the same time; Niu et al. (2023) removes partial noisy samples with large gradient and encourages model weights to go to a flat minimum to achieve stable TTA. In our work, we combine test-time adaptation with contrastive learning, in order to further facilitate domain adaptation in out-of-context news detection.
3. Problem Statement
@@
In out-of-context news detection, given the news image-text pair $(x_{img},x_{txt})$, the detection model is expected to predict whether it is falsified, i.e. out-of-context. If the image and text are not from the same news post, this image-text pair is considered as falsified (labeled as True), otherwise not falsified (labeled as False).
@@
In this work, we tackle the challenge of domain adaptive out-of-context news detection. During training, we have access to the labeled source domain data and unlabeled target domain data. During testing, we evaluate our model’s performance on the labeled target domain data. The source domain and target domain are mutually exclusive. The target domain training data and testing data are mutually exclusive as well. Concretely, the dataset is comprised of news from $N$ domains. Each time we select $M$ domains as the target domain, $\mathcal{D}^{T}=\{\mathcal{D}^{T_{m}}\}_{m=1}^{M}$, and the rest $N-M$ domains as the source domain, $\mathcal{D}^{S}=\{\mathcal{D}^{S_{n}}\}_{n=1}^{N-M}$. The $n$-th source domain $\mathcal{D}^{S_{n}}$ contains labeled data, $\mathcal{D}^{S_{n}}=\{(x_{i}^{S_{n}},y_{i}^{S_{n}})\}_{i=1}^{|S_{n}|}$, where $x^{S_{n}}_{i}=(x_{img},x_{txt})_{i}^{S_{n}}$ denotes the $i$-th news image-text pair within domain $\mathcal{D}^{S_{n}}$, and $y_{i}^{S_{n}}$ denotes whether this image-text pair is falsified or not (True or False). In contrast, the $m$-th target domain contains unlabeled data: $\mathcal{D}^{T_{m}}=\{x_{j}^{T_{m}}\}_{j=1}^{|T_{m}|}$, where $x_{j}^{T_{m}}=(x_{img},x_{txt})_{j}^{T_{m}}$.
@@
4. Our Approach: ConDA-TTA
In this section, we describe our proposed approach for domain adaptive out-of-context news detection: Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA). The overall architecture of ConDA-TTA is illustrated in Figure 2. It consists of three components: (i) Multimodal Feature Encoder, which encodes the image and text pair into a multimodal representation; (ii) Domain-Invariant Feature Learning, which applies contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature; and (iii) Test-Time Adaptation, which leverages unlabeled test set statistics to further adapt the model to the target domain. Finally, we discuss the training details of our approach.
Figure 2. The overall model architecture of ConDA-TTA. We first use the (i) Multimodal Feature Encoder to encode out-of-context news and its augmentation into multimodal representations. Then in the (ii) Domain-Invariant Feature Learning, we apply contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature shared by different domains. Finally, we adopt the (iii) Test-Time Adaptation to update statistic-related model parameters in the evaluation phase.
4.1. Multimodal Feature Encoder
In order to capture the semantic nuance between original news and out-of-context news, we use MLLM to directly encode the news text and image into a meaningful multimodal representation.
@@ 4.1. Multimodal Feature Encoder
(1) $\textbf{x}^{S}=\mathrm{MLLM}((x_{img},x_{txt})^{S}),\quad\textbf{x}^{T}=\mathrm{MLLM}((x_{img},x_{txt})^{T}),$
@@
where $\textbf{x}^{S}$ and $\textbf{x}^{T}$ respectively denote the multimodal feature representation of news image-text pair $(x_{img},x_{txt})$ from the source and target domain.
4.2. Domain-Invariant Feature Learning
@@
After encoding the news, we introduce a domain-invariant feature learning module to learn the domain-invariant feature from the news’ multimodal representation. Specifically, this module adopts contrastive learning to learn a more separable representation space and adopts MMD to capture the invariant features among different domains.
@@
We first use a projection head to project the multimodal feature representation into a lower dimensional space. The projection head is composed of two linear layers, and can be expressed as follows:
(2) $\textbf{z}^{S}=\mathrm{Projection}(\textbf{x}^{S}),\quad\textbf{z}^{T}=\mathrm{Projection}(\textbf{x}^{T}).$
@@
Next, we aim to train the projection head so that it can capture the domain-invariant features from $\textbf{x}^{S}$ and $\textbf{x}^{T}$. This is enabled by applying contrastive learning and the maximum mean discrepancy. Note, no label information is required within this module.
@@
Contrastive Learning: Inspired by (Bhattacharjee et al., 2023), we leverage contrastive learning to learn more separable representations of the input multimodal features in the projected space. In this way, we expect that the learned representations of similar semantic meanings stay closer, and the learned representations of dissimilar semantic meanings stay further apart.
@@
To achieve this, we adopt the contrastive loss used in  (Bhattacharjee et al., 2023). Applying contrastive loss pulls closer the positive pairs and pushes further apart the negative pairs. We augment each news item in order to construct the positive and negative pairs later. For each news item $\textbf{x}_{i}^{S}$ in a source domain training batch $\{\textbf{x}_{i}^{S}\}_{i=1}^{b}$, we generate one augmentation of it and encode it by the same multimodal feature encoder to obtain its representation, denoted as $\textbf{x}_{i+}^{S}$. In this way, we obtain $2|b|$ news items in the augmented batch. After projecting the input multimodal feature to the lower dimensional space, the positive pair is formed by the data item and its corresponding augmentation item, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{i+}^{S})$; the negative pairs are formed by the data item and the rest of the $2(|b|-1)$ items within the augmented batch, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})$. The contrastive loss is expressed as follows:
(3) $\mathcal{L}_{ctr}^{S}=-\sum_{(i,i_{+})\in b}\log\frac{\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{i_{+}}^{S})/t)}{\sum_{k=1,k\neq i}^{2|b|}\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})/t)}.$
@@
Here, $b$ denotes the current batch and $|b|$ denotes the size of the current batch. $S$ denotes the source domain. $\textbf{z}_{i}^{S}$ and $\textbf{z}_{j}^{S}$ denote the learned representations of the $i$-th data item and its augmentation in the projected space. $t$ denotes the temperature coefficient. $\mathrm{sim}(\cdot)$ denotes the similarity metric. Here, cosine similarity is being used. The data augmentation on the target domain training batch and the contrastive loss computation for the target domain $\mathcal{L}_{ctr}^{T}$ is applied in the same way.
The performance of contrastive learning is largely influenced by the transformation used to generate positive samples (Tian et al., 2020). Following the data augmentation experiments in (Bhattacharjee et al., 2023), we include text-level: {synonym replacement, random swap, random crop}, and image-level: {random resize and crop, random horizontal flip, Gaussian blur} into the set of candidate data augmentation techniques (detailed descriptions can be found in the supplementary material). We also include the combination of text-level and image-level augmentations as the candidates. Our experiments show that all the above strategies result in similar performances (in terms of accuracy on the target domain), but applying Gaussian blur to the image results in the best overall performance (averaged over all domain adaptation settings) on the target domain. Thus, in our approach, we apply Gaussian blur to the image while keeping the text unaltered to generate the positive sample of the anchor item. Note, the augmentation preserves the label.
@@
Maximum Mean Discrepancy (MMD): Now that we have learned more separable representations, $\textbf{z}^{S}$ and $\textbf{z}^{T}$, in the projected space, we want to further ensure that these representations capture domain-invariant feature. To achieve this, we adopt maximum mean discrepancy (MMD) (Gretton et al., 2012). MMD is a kernel-based statistical test used to determine if two samples are drawn from different distributions (Gretton et al., 2012). We use it as a loss function to measure the discrepancy between two distributions. Here, we regard the learned representations $\textbf{z}^{S}$ and $\textbf{z}^{T}$ as the samples from two random variables $Z^{S}$ and $Z^{T}$, and we want their probability distributions to have as small as possible discrepancy. In this way, the distribution of $Z^{S}$ and $Z^{T}$ conditions less on a specific domain. Alternatively speaking, if MMD is small enough, we regard that it removes the domain-specific features from z and keeps the domain-invariant features.
@@
(4) $\mathrm{MMD}^{2}[\mathcal{F},Z^{S},Z^{T}]$ $=\left[\sup_{||f||_{\mathcal{H}}\leq 1}(\mathbf{E}[f(Z^{S})]-\mathbf{E}[f(Z^{T})])\right]^{2}$
@@
Following prior work (Pan et al., 2010; Long et al., 2015; Bhattacharjee et al., 2023), we compute the MMD between features in the lower dimensional space (i.e. the projected z space). In our implementation, we use the empirical MMD (Gretton et al., 2012), which is expressed as follows:
(5) $\mathrm{MMD}[\mathcal{F},Z^{S},Z^{T}]=$ $\left[\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{S})-\frac{2}{mn}\sum_{i,j=1}^{m,n}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{T})\right.$
@@
(6) $k(\textbf{z}_{i},\textbf{z}_{j})=\exp\left(-\frac{||\textbf{z}_{i}-\textbf{z}_{j}||^{2}_{2}}{2\sigma^{2}}\right).$

### v2

@@ Abstract
Out-of-context news is a common type of misinformation on online media platforms. This involves posting a caption, alongside a mismatched news image. Existing out-of-context news detection models only consider the scenario where pre-labeled data is available for each domain, failing to address the out-of-context news detection on unlabeled domains (e.g. news topics or agencies). In this work, we therefore focus on domain adaptive out-of-context news detection. In order to effectively adapt the detection model to unlabeled news topics or agencies, we propose ConDA-TTA (Contrastive Domain Adaptation with Test-Time Adaptation) which applies contrastive learning and maximum mean discrepancy (MMD) to learn domain-invariant features. In addition, we leverage test-time target domain statistics to further assist domain adaptation. Experimental results show that our approach outperforms baselines in most domain adaptation settings on two public datasets, by as much as 2.93% in F1 and 2.08% in accuracy.
Learning Domain-Invariant Features for Out-of-Context News Detection
Yimeng Gu1, Mengqi Zhang2, Ignacio Castro1, Shu Wu3, Gareth Tyson4 1Queen Mary University of London, 2Shandong University, 3CRIPC, Chinese Academy of Science, 4The Hong Kong University of Science and Technology (GZ) yimeng.gu@qmul.ac.uk
1 Introduction
@@
Online news platforms suffer from the release and spread of misinformation. A common Fazio (2020) yet subtle type of misinformation is image repurposing, also called out-of-context news. Unlike DeepFake (Dolhansky et al., 2020) models which generate non-existing images, this approach attaches a false claim with a real image outside of its original context (see Figure 1). This lowers the technical threshold Fazio (2020) of producing misinformation since anyone can trivially attach a fabricated caption to an image and post it online, thereby enlarging its scope of occurrence.
@@
Figure 1: Examples of out-of-context news of three different news topics from the Twitter-COMMs dataset.
@@
Out-of-context news detection has received a growing attention in recent years. Existing works (Biamby et al., 2022; Luo et al., 2021; Abdelnabi et al., 2022; Mu et al., 2023) fine-tune pre-trained vision and text encoders to represent the image and text. More recent works Dai et al. (2024); Zhang et al. (2023) query Multimodal Large Language Models (MLLMs) to obtain the prediction with explanations. However, their approaches do not address the model adaptation to entirely new topics or news agencies. Due to the wide array of news topics shared online, it is infeasible to re-train the detector every time a new topic emerges. Considering annotation costs, it is equally critical to build a detection model that can adapt to other news agencies with minimal efforts.
@@
Unfortunately, prior works neglect addressing the domain adaption issue on out-of-context news detection. They tend to capture domain-specific knowledge not shared across different domains. Although this knowledge is helpful for improving the model’s performance on annotated domains, it results in inferior performance on unannotated domains Wang et al. (2018). In light of this, we believe that learning domain-invariant features will help the detection model adapt to unlabeled domains. However, this comes with several challenges. First, existing out-of-context news datasets only cover a small number of topics/agencies. This makes it challenging for the detection model to learn domain-invariant features, because the learned features might be biased towards the limited number of source domains. Second, in (social media) news dataset, the data distributions of different domains vary, since news from different domains tend to have different writing styles and image styles. This makes domain-invariant knowledge learning challenging. It is critical to ensure that transferable domain-invariant knowledge is actually learned. Otherwise, the model may bypass learning it and end up learning semantic patterns as a shortcut to fulfill the classification task (Li et al., 2023b; Chi et al., 2023).
@@
To tackle the aforementioned challenges, we propose Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA) for domain adaptive out-of-context news detection. ConDA-TTA first uses an MLLM to directly encode the image and text into a multimodal feature representation. After that, we adopt the contrastive loss to learn a representation in the projected space where the original news stays further away from the out-of-context news. To overcome the shortage of labeled source domains, we take advantage of the unlabeled target domain data as well, to learn less biased domain-invariant features. To ensure that domain-invariant features are captured from the learned representations, we then apply Maximum Mean Discrepancy (MMD) to reduce the discrepancy of the learned representations between the source and target domain. Finally, we incorporate Test-Time Adaptation (TTA) to update the statistic-related model parameters in the evaluation phase so that the model can better adapt to the target domain.
@@
  • We are the first to investigate domain adaptation in out-of-context news detection. To address the challenges, we propose a novel approach that learns domain-invariant features through MMD and TTA.
@@
  • We evaluate our approach and demonstrate its effectiveness. Notably, it outperforms the baseline by as much as 2.93% in F1 when the domain is defined as news topic. Additionally, it outperforms the baseline by as much as 1.82% in F1 when the domain is defined as news agency.
@@
  • We conduct a comprehensive ablation study and show that MMD is the most contributing component to the domain adaptation on Twitter-COMMs, while TTA is the most contributing component on NewsCLIPpings.
2 Related Work
2.1 Out-of-Context News Detection
@@
Prior works on out-of-context news detection mainly adopt three technical routes: (i) fine-tuning or (ii) prompting large vision and language models, and (iii) leveraging synthetic information. Biamby et al. (2022) fine-tunes CLIP (Radford et al., 2021) to detect out-of-context news. Similarly, Luo et al. (2021) fine-tunes both CLIP and VisualBert (Li et al., 2019). Abdelnabi et al. (2022) additionally leverages the retrieved textual and visual evidences. These works involve fine-tuning large multimodal pre-trained models, which is very compute-intensive and does not enable domain adaptation.
@@
Later work leverages synthetic multimodal information. Shalabi et al. (2023) generates augmented image from text, and generates augmented text from image. Yuan et al. (2023) extracts stances from external multimodal evidences to enhance the detection. More recent works focus on enhancing interpretability. Qi et al. (2024) adopts two-stage instruction tuning on Instruct-BLIP Dai et al. (2024) to provide accurate and persuasive explanations to the prediction. Zhang et al. (2023) uses neural symbolic model to enhance model interpretability. Shalabi et al. (2024) fine-tunes MiniGPT-4 to detect out-of-context news. Although effective, the above works do not address the domain adaptation problem in out-of-context news detection.
2.2 Domain Adaptive Fake News Detection
@@
To the best of our knowledge, no prior work has explored domain adaptive out-of-context news detection. However, existing works have investigated domain adaptation in fake news detection — a task close to out-of-context news detection. Many works Wang et al. (2018); Zhang et al. (2020, 2021); Yuan et al. (2021) adopt adversarial learning to learn the domain-invariant feature. Furthermore, by incorporating user comments and user-news interaction information, Mosallanezhad et al. (2022) adopts reinforcement learning that exploits cross-domain and within-domain knowledge to achieve robustness in the target domain. Yue et al. (2022) proposes a contrastive domain adaptation method in order to reduce the intra-class discrepancy and enlarge the inter-class discrepancy. Lin et al. (2023) employs prompt engineering to learn language-agnostic contextual representations and models the domain-invariant structural features from the propagation threads.
In contrast, our work focuses on the domain adaptation in out-of-context news detection and adopts both contrastive learning and test-time adaptation to address the unique challenges (Section  1) of this task.
3 Problem Statement
@@
In out-of-context news detection, given the news image-text pair $(x_{img},x_{txt})$, the detection model is expected to predict whether it is falsified, i.e. out-of-context. If the image and text are not from the same news post, this news is considered as falsified (labeled as True), otherwise not falsified (labeled as False).
@@
In this work, we tackle the challenges of domain adaptive out-of-context news detection. During training, we have access to the labeled source domain data and unlabeled target domain data. During testing, we evaluate our model’s performance on the labeled target domain data. The source domain and target domain are mutually exclusive. Concretely, the dataset is comprised of news from $N$ domains. Each time we select $M$ domains as the target domain, $\mathcal{D}^{T}=\{\mathcal{D}^{T_{m}}\}_{m=1}^{M}$, and the rest $N-M$ domains as the source domain, $\mathcal{D}^{S}=\{\mathcal{D}^{S_{n}}\}_{n=1}^{N-M}$. The $n$-th source domain $\mathcal{D}^{S_{n}}$ contains labeled data, $\mathcal{D}^{S_{n}}=\{(x_{i}^{S_{n}},y_{i}^{S_{n}})\}_{i=1}^{|S_{n}|}$, where $x^{S_{n}}_{i}=(x_{img},x_{txt})_{i}^{S_{n}}$ denotes the $i$-th news image-text pair within domain $\mathcal{D}^{S_{n}}$, and $y_{i}^{S_{n}}$ denotes whether this image-text pair is falsified or not. In contrast, the $m$-th target domain contains unlabeled data: $\mathcal{D}^{T_{m}}=\{x_{j}^{T_{m}}\}_{j=1}^{|T_{m}|}$, where $x_{j}^{T_{m}}=(x_{img},x_{txt})_{j}^{T_{m}}$.
@@
4 Our Approach: ConDA-TTA
In this section, we describe our proposed approach for domain adaptive out-of-context news detection: Contrastive Domain Adaptation with Test-Time Adaptation (ConDA-TTA). The overall architecture of ConDA-TTA is illustrated in Figure 2. It consists of three components: (i) Multimodal Feature Encoder, which encodes the image and text pair into a multimodal representation; (ii) Contrastive Domain Adaptation, which applies contrastive learning to learn a more separable representation space and maximum mean discrepancy (MMD) to learn the domain-invariant feature; and (iii) Test-Time Adaptation, which leverages unlabeled test set statistics to further adapt the model to the target domain.
Figure 2: The model architecture of ConDA-TTA. We first use the (i) Multimodal Feature Encoder to encode the news and its augmentation into multimodal representations. Then in the (ii) Contrastive Domain Adpatation, we apply contrastive learning and maximum mean discrepancy (MMD) to learn the domain-invariant feature. Finally, we adopt the (iii) Test-Time Adaptation to update statistic-related model parameters in the evaluation phase.
4.1 Multimodal Feature Encoder
In order to capture the semantic differences between original news and out-of-context news, we use MLLM to directly encode the news text and image into a meaningful multimodal representation.
@@ 4.1 Multimodal Feature Encoder
$\textbf{x}^{S}$ $=\mathrm{MLLM}((x_{img},x_{txt})^{S}),$ (1)
$\textbf{x}^{T}$ $=\mathrm{MLLM}((x_{img},x_{txt})^{T}),$
@@
where $\textbf{x}^{S}$ and $\textbf{x}^{T}$ respectively denote the multimodal feature representation of the news image-text pair $(x_{img},x_{txt})$ from the source and target domain.
4.2 Contrastive Domain Adaptation
@@
After encoding the news, we introduce the contrastive domain adaptation module. Specifically, this module first adopts contrastive learning to learn a more separable representation space for news image-text pairs and then uses MMD to capture the invariant features among different domains.
@@
We first use a projection head to project the multimodal feature representation into a lower dimensional space. The projection head can be expressed as follows:
$\textbf{z}^{S}=\mathrm{Projection}(\textbf{x}^{S}),\quad\textbf{z}^{T}=\mathrm{Projection}(\textbf{x}^{T}).$ (2)
@@
Next, we aim to train the projection head so that it can learn a more separable representation space and capture the domain-invariant features from $\textbf{x}^{S}$ and $\textbf{x}^{T}$ afterwards. This is achieved by applying the contrastive learning and maximum mean discrepancy respectively. Note, no label information is required within this module.
@@
Contrastive Learning: Contrastive learning has been shown effective in learning better representations for the classification task Qian et al. (2022). Inspired by Bhattacharjee et al. (2023), we leverage contrastive learning to learn more separable representations of the input multimodal features in the projected space. In this way, we expect that the learned representations of similar semantic meanings stay closer, and the learned representations of dissimilar semantic meanings stay further apart, so that it could make the classification easier.
@@
To achieve this, we adopt the contrastive loss used in  Bhattacharjee et al. (2023). Applying contrastive loss pulls closer the positive pairs and pushes further apart the negative pairs. To construct the positive and negative pairs, we first augment each news item. For each news item $\textbf{x}_{i}^{S}$ in a source domain training batch $\{\textbf{x}_{i}^{S}\}_{i=1}^{b}$, we generate one augmentation of it and encode it by the same multimodal feature encoder to obtain its representation, denoted as $\textbf{x}_{i+}^{S}$. In this way, we obtain $2|b|$ news items in the augmented batch. After projecting the input multimodal feature to the lower dimensional space, the positive pair is formed by the data item and its corresponding augmentation item, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{i+}^{S})$; the negative pairs are formed by the data item and the rest of the $2(|b|-1)$ items within the augmented batch, denoted as $(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})$. The contrastive loss is expressed as follows:
$\mathcal{L}_{ctr}^{S}=-\sum_{(i,i_{+})\in b}\log\frac{\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{i_{+}}^{S})/t)}{\sum_{k=1,k\neq i}^{2|b|}\exp(\mathrm{sim}(\textbf{z}_{i}^{S},\textbf{z}_{k}^{S})/t)}.$ (3)
@@
Here, $b$ denotes the current batch and $|b|$ denotes the size of the current batch. $S$ denotes the source domain. $\textbf{z}_{i}^{S}$ and $\textbf{z}_{j}^{S}$ denote the learned representations of the $i$-th data item and its augmentation in the projected space. $t$ denotes the temperature coefficient. $\mathrm{sim}(\cdot)$ denotes the similarity metric. Here, cosine similarity is being used. The data augmentation on the target domain training batch and the contrastive loss computation for the target domain $\mathcal{L}_{ctr}^{T}$ are applied in the same way. Note, the augmentation preserves the label.
@@
Maximum Mean Discrepancy (MMD): Now that we have learned more separable representations, $\textbf{z}^{S}$ and $\textbf{z}^{T}$, in the projected space, we want to further capture domain-invariant features. To achieve this, we adopt maximum mean discrepancy (MMD) Gretton et al. (2012). MMD is a kernel-based statistical test used to determine if two samples are drawn from different distributions Gretton et al. (2012). We use it as a loss function to measure the discrepancy between two distributions. Here, we regard the learned representations $\textbf{z}^{S}$ and $\textbf{z}^{T}$ as the samples from two random variables $Z^{S}$ and $Z^{T}$, and we want their probability distributions to have as small as possible discrepancy. In this way, the distribution of $Z^{S}$ and $Z^{T}$ conditions less on a specific domain. Alternatively speaking, if MMD is small enough, we regard that it removes the domain-specific features from z and keeps the domain-invariant features.
@@
$\mathrm{MMD}^{2}$ $=\left[\sup_{||f||_{\mathcal{H}}\leq 1}(\mathbf{E}[f(Z^{S})]-\mathbf{E}[f(Z^{T})])\right]^{2}$ (4)
@@
Following prior work (Pan et al., 2010; Long et al., 2015; Bhattacharjee et al., 2023), we compute the MMD of features in the lower dimensional space (i.e. the projected z space). In our implementation, we use the empirical MMD Gretton et al. (2012), which is expressed as follows:
$\mathrm{MMD}=$ $\left[\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{S})-\frac{2}{mn}\sum_{i,j=1}^{m,n}k(\textbf{z}_{i}^{S},\textbf{z}_{j}^{T})\right.$ (5)
@@
$k(\textbf{z}_{i},\textbf{z}_{j})=\exp\left(-\frac{||\textbf{z}_{i}-\textbf{z}_{j}||^{2}_{2}}{2\sigma^{2}}\right).$ (6)